-- PaulSmith? - 24 Feb 2007

SDT? Observer1: Nicolas Bouche, Peter Buschkamp Observer2: Paul Smith Inst. Support: TOs: Ed Lundin and David Gonzalez

1730-1800: biases and darks.

1827 Open dome to do flats Flats done (PA=180 deg) rVgB , (PA=0 deg) rV

Clear at start, but a bright moon right overhead (slightly more than first quarter).

2000: Needing to spend quite a lot of time to get the collimation close by hand so that the automatic routine will work. Were having trouble getting our directions right! We learn from Olga that we need to zero coefficients that we don'r want to make changes to before "sending" the coefficients that we do want to change. This requirement was not clear to either the observers nor the operators.

2020: Finally sort it out. Do OB=L95_gr; seeing 1.0". OB=L95_BV; seeing 0.7".

2030: To M81 and WF, the OB=KochanekM81V; seeing 0.8-0.9".

2050: Move to Will1_1_R, do WF, then OB=Will1_1_R; seeing 0.7". WF then OB=Will1_2_R; seeing 0.6-0.8".

2137: Move to cosmos_rSLOAN_3 (LBC gui freezes - restart; problems getting LBC up again; CCD controller error). Call John Hill. Completely power down LBC and then reboot. LBC finally comes up correctly

2220: Do WF, then OB=cosmos_rSLOAN3; seeing 0.7-0.8". At start of every integration we get a "left main mirror fails to focus" error. Do WF, then OB=cosmos_rSLOAN_4; seeing 0.8". A fair amount of guiding errors are seen in these long (360s) exposures resulting in slightly elongated PSFs. Things could probably be greatly improved by guiding at a much faster rate.

2348: Do WF, then OB=cosmos_rSLOAN5; seeing 0.7-1.0". Do WF then OB=cosmos_rSLOAN6; seeing 0.8". Do WF then OB=cosmos_rSLOAN7; seeing 0.6-0.7" (some excellent pictures here; PSFs rounder than usual).

0153: Move to standard star field (trouble with LBC gui again). Need to spend some time recovering from the LBC "freeze up"

0210: Do WF then OB=PG1047_BV and then OB=PG1047_gr.

0220: Move to fuobj8_2_r, do WF and then OB=fuobj8_2_r; seeing 0.8-0.9". Moon setting.

0250: Move to Lockman hole, do WF, then OB=Egami_Lockman_XMM_B_2; seeing 0.8-1.0". WF then OB=Egami_Lockman_XMM_B_1 (this OB is a repeat of a couple of nights ago when the image quality was not very good); seeing 0.8-1.0"

0417: Move to M53, do WF then OB=N5024_short; seeing 0.9-1.0". Some really nice images here. Do OB=N5024_deep; seeing 0.9". WF the OB=N5024_deep; seeing 0.8-0.9".

0515: Move to standard star field, do WF and then OB=PG1633_U, PG1633_BV,PG1633_gr.

0600: Do morning flats. UBVgr (PA = 0 deg); U (PA = 180 deg).

0635: End of night, end of SDT observing run. Pretty good final night. A few technical glitches, but nothing too serious. The SDT observers would like the thank the LBTO staff for an enjoyable and productive stay at the MGIO.
Topic revision: r5 - 25 Feb 2007, PaulSmith
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback